OpenStudy (anonymous):

Help medal and fan!! 9th grade agebra1

OpenStudy (yanasidlinskiy):

Can you please post 1 question at a time?

OpenStudy (anonymous):

sure @YanaSidlinskiy

OpenStudy (anonymous):

) For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. (hint: you need to find the “k” value) X Y 6 9 8 10 10 11 and ok @ryan>>>

OpenStudy (yanasidlinskiy):

Like....Go ahead and close it and post 1 at a time and work your way through!:) I'll be glad to help you!:) I will let others go ahead and help you!:) But if it confuses...just tag me and I'll help!:)

OpenStudy (anonymous):

1) For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. (hint: you need to find the “k” value) X Y 6 9 8 10 10 11

OpenStudy (anonymous):

@ryan>>> can u help plz

OpenStudy (anonymous):

@YanaSidlinskiy they aren't helping

OpenStudy (yanasidlinskiy):

In order for y to vary directly with x, it must satisfy the equation y = kx. So, for the values, plug them in to see: 6=k(9) 8 =k(10) 10= k(11)

OpenStudy (anonymous):

2) For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. (Hint: you need to find the “k” value) X Y 8 6 12 9 20 15

OpenStudy (anonymous):

the line is straight @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Same here. Do the same like I did. Check this out! This might help you out a little....If not just tag me again.....But I think you should understand it without me explaining it to you:) https://answers.yahoo.com/question/index?qid=20131111193024AAwLHMT